# Label Distribution Learning using the Squared Neural Family on the Probability Simplex

### v1

@@ 5 SNEFY-LDL
The proof is provided in the Appendix. With the closed-form kernel function in Eq. (10), we can construct the conditional SNEFY distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ in the form of Eq. (7). This model provides us with the freedom to choose any sufficient statistic $\bm{t}_{2}(\cdot)$ that is used to transform the input sample $\bm{x}$ from the original $d$-dimensional space to the latent $D_{2}$-dimensional space. To capture the non-linearity between input samples and their label distributions, deep neural networks can be leveraged to construct $\bm{t}_{2}(\cdot)$. The input can also be extended beyond the vector-format samples to data with special structures, like images, texts and graphs, where we can respectively leverage Convolutional Neural Networks (CNNs) (Venkatesan and Li, 2017), Transformers (Vaswani et al., 2017) and Graph Neural Networks (GNNs) (Kipf and Welling, 2017) to construct $\bm{t}_{2}(\cdot)$ for end-to-end learning.
The conditional distribution formulation $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ also provides a closed form of mean, variance and co-variance of the label distribution probabilities conditioned on the input sample $\bm{x}$. About this, we have the following theorem:
@@
Assuming the label distribution vector $\bm{\ell}$ follows the SNEFY conditional distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ in Eq. (7) with the kernel function $\bm{k}_{\sigma,\bm{t}_{1},\bm{t}_{2},\mu_{1}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})$ given in Eq. (10), under the setting that $\bm{t}_{1}(\bm{\ell})=(\log\bm{\ell}^{y_{1}},\log\bm{\ell}^{y_{2}},\cdots,\log\bm{\ell}^{y_{L}})$, $\sigma=\exp$, and $\mu_{1}(d\bm{\ell})=d\bm{\ell}$, as well as the constraint that $\bm{W}_{1}>-1/2$ elementwise, for the affiliation probability to the $r$th label $y_{r}\in\mathcal{Y}$, $\bm{\ell}^{y_{r}}$, we have its conditional mean $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ as
@@
The proof is provided in the Appendix. Given the fitted distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$, the mean $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ can be used to predict the unknown label distribution as the expectation over all values in the simplex. We can use the variance $\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]$ to quantify label distribution prediction uncertainties. We can also use $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ with $\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]$ to construct confidence intervals for label distribution predictions by applying Chebyshev’s inequality (Grimmett and Stirzaker, 2020). The covariance $\mathrm{Cov}[\bm{\ell}^{y_{r}},\bm{\ell}^{y_{s}}|\bm{x}]$ is helpful for us to understand the correlations between two different labels. More importantly, all the statistics are conditioned on the input sample $\bm{x}$, guiding us to make instance-wise inferences.
@@
There are numerous metrics to measure the consistency between two label distributions, like Chebyshev distance, Kullback-Leibler divergence and Cosine coefficient (Geng, 2016). Instead of optimizing these metrics, the MLE based learning objective in Eq. (17) provides an alternative way to train the LDL model. The fitted distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ can be applied to various downstream tasks, and its closed-form mean is able to offer accurate label distribution predictions that are generic to varying evaluation metrics.
Algorithm Description and Time Complexity
We train the SNEFY-LDL model with stochastic gradient descent. The training procedure is shown in Algorithm 1. The model parameters $\bm{V}$, $\bm{\Theta}$ and $\bm{t}_{2}$ are first initialized by random numbers. We then iteratively select a batch of training samples, calculate the batched likelihoods with Eq. (7), and update parameters $\bm{V}$, $\bm{\Theta}$ and $\bm{t}_{2}$ by descending them along the gradient of batched negative log likelihoods. Taking the epoch number as a constant and assuming the latent layers of $\bm{t}_{2}$ have neurons in the same scale as the neuron number in the last layer $D_{2}$, the time complexity of Algorithm 1 is $\mathcal{O}(N(mn^{2}+Ln^{2}+dD_{2}+D_{2}^{2}))$, which is linear to the number of training samples $N$, making the algorithm able to scale to large datasets. For any sample $\bm{x}$ with an unknown label distribution, its label distribution can be predicted fast in time complexity $\mathcal{O}(mn^{2}+Ln^{2}+dD_{2}+D_{2}^{2})$ using the closed-form mean formulation in Eq. (11).
@@ 6 Experiments
In this section, extensive experiments are carried out to evaluate SNEFY-LDL’s performance in label distribution prediction and prediction uncertainty quantification.
Benchmark Datasets
@@
We use fourteen datasets (Geng, 2016) to benchmark our experiments, including the Movie dataset containing label distributions on five movie rating scales (from one to five), the Natural Scene dataset with label distributions constructed by inconsistent multi-label ranking on natural scene images, the facial expression datasets SBU_3DFE and SJAFFE with label distributions on six emotions, as well as the ten Yeast series datasets built from ten biological experiments on a budding yeast, where labels are defined by different time points and label distributions are constructed by different gene expression levels at varying time points. The statistics of the fourteen benchmark datasets are summarized in Table 1.
Baseline Methods
We compare the proposed SNEFY-LDL model with eight state-of-the-art LDL baselines:
  • PT-SVM (Geng, 2016) transforms LDL into a multi-class classification problem and uses SVM as the classifier.
  • AA-KNN (Geng, 2016) adapts KNN from the single-label classification setting to the LDL setting.
  • CPNN (Geng et al., 2013) uses a neural network to model the feature-label joint distribution.
  • SA-BFGS (Geng, 2016) applies the maximum entropy model to characterize label distributions.
  • LDLF (Shen et al., 2017) employs the differentiable decision trees to model label distributions.
  • DF-LDL (González et al., 2021a) learns label distributions in a decomposition-fusion manner.
  • LDL-SCL (Jia et al., 2021) considers the label distribution correlations between similar samples.
  • LDL-LRR (Jia et al., 2023a) preserves the label importance ranking in label distribution modeling.
No. Dataset #Examples #Features #Labels
1 Movie 7,755 1,869 5
2 Natural Scene 2,000 294 9
3 SBU_3DFE 2,500 243 6
4 SJAFFE 213 243 6
5 Yeast_alpha 2,465 24 18
6 Yeast_cdc 2,465 24 15
7 Yeast_cold 2,465 24 14
8 Yeast_diau 2,465 24 7
9 Yeast_dtt 2,465 24 6
10 Yeast_elu 2,465 24 6
11 Yeast_heat 2,465 24 4
12 Yeast_spo 2,465 24 4
13 Yeast_spo5 2,465 24 3
14 Yeast_spoem 2,465 24 2
Table 1: Summary of the fourteen benchmark datasets.
Experimental Settings
Following (Geng, 2016), we use ten-fold cross-validation to evaluate the label distribution prediction performance, by randomly splitting each dataset into ten chunks, and iteratively choosing a chunk as the test set and taking the remaining nine chunks as the training set. LDL models are then trained on the training set and evaluated on the test set for ten rounds. Averaged performance is used to compare SNEFY-LDL with baseline methods.
@@ 14 Yeast_spoem 2,465 24 2
The closed-form mean in Eq. (11) is used to make comparisons on the label distribution prediction task with the baselines which all return point estimates on the simplex. However, we note that SNEFY-LDL is a full conditional density model and its utility is broader than the baselines, as we illustrate in the active learning and ensemble learning experiments. When implementing SNEFY-LDL, $m$ and $n$ are respectively set to 32 and 64, $D_{2}$ is set as equal to $n$ and a one-layer neural network with ReLU activation is used to construct $\bm{t}_{2}$. The model is trained for 100 epochs with batch size 16. Weight clipping (Arjovsky et al., 2017) is used to control $\bm{W}_{1}>-1/2$ after each parameter update. For the eight baseline models, we use the implementations in the PyLDL library [^1] https://github.com/SpriteMisaka/PyLDL and keep the default hyperparameters.
Dataset PT-SVM AA-KNN CPNN SA-BFGS LDLF DF-LDL LDL-SCL LDL-LRR SNEFY-LDL $\Delta$
Movie 0.9332 $\bullet$ 0.9279 $\bullet$ 0.9244 $\bullet$ 0.9339 $\bullet$ 0.8596 $\bullet$ 0.9357 $\spadesuit$ 0.9036 $\bullet$ 0.9353 $\spadesuit$ 0.9354 -0.0002
Natural_Scene 0.7404 $\spadesuit$ 0.7112 $\bullet$ 0.6249 $\bullet$ 0.7168 $\bullet$ 0.7304 $\spadesuit$ 0.7377 $\spadesuit$ 0.6716 $\bullet$ 0.7455 $\spadesuit$ 0.7524 +0.0068
SBU_3DFE 0.9244 $\bullet$ 0.9215 $\bullet$ 0.9169 $\bullet$ 0.9383 $\bullet$ 0.9359 $\bullet$ 0.9386 $\bullet$ 0.9249 $\bullet$ 0.9284 $\bullet$ 0.9427 +0.0042
SJAFFE 0.9336 $\bullet$ 0.9490 $\spadesuit$ 0.9302 $\bullet$ 0.9543 $\spadesuit$ 0.9593 $\circ$ 0.9646 $\circ$ 0.9604 $\circ$ 0.9635 $\circ$ 0.9507 -0.0140
Yeast_alpha 0.9941 $\bullet$ 0.9938 $\bullet$ 0.9940 $\bullet$ 0.9946 $\circ$ 0.9943 $\spadesuit$ 0.9946 $\circ$ 0.9945 $\circ$ 0.9946 $\circ$ 0.9943 -0.0004
Yeast_cdc 0.9928 $\bullet$ 0.9924 $\bullet$ 0.9927 $\bullet$ 0.9933 $\circ$ 0.9932 $\circ$ 0.9933 $\circ$ 0.9932 $\circ$ 0.9933 $\circ$ 0.9929 -0.0004
Yeast_cold 0.9879 $\bullet$ 0.9872 $\bullet$ 0.9875 $\bullet$ 0.9885 $\circ$ 0.9876 $\bullet$ 0.9884 $\spadesuit$ 0.9871 $\bullet$ 0.9885 $\circ$ 0.9883 -0.0002
Yeast_diau 0.9870 $\bullet$ 0.9867 $\bullet$ 0.9860 $\bullet$ 0.9879 $\circ$ 0.9873 $\spadesuit$ 0.9879 $\circ$ 0.9849 $\bullet$ 0.9879 $\circ$ 0.9873 -0.0006
Yeast_dtt 0.9934 $\bullet$ 0.9934 $\bullet$ 0.9938 $\spadesuit$ 0.9941 $\circ$ 0.9936 $\spadesuit$ 0.9939 $\spadesuit$ 0.9931 $\spadesuit$ 0.9941 $\circ$ 0.9939 -0.0002
Yeast_elu 0.9933 $\bullet$ 0.9932 $\bullet$ 0.9936 $\spadesuit$ 0.9940 $\circ$ 0.9939 $\circ$ 0.9941 $\circ$ 0.9939 $\circ$ 0.9940 $\circ$ 0.9936 -0.0005
Yeast_heat 0.9871 $\bullet$ 0.9867 $\bullet$ 0.9874 $\bullet$ 0.9880 $\circ$ 0.9877 $\spadesuit$ 0.9881 $\circ$ 0.9873 $\spadesuit$ 0.9880 $\circ$ 0.9878 -0.0003
Yeast_spo 0.9765 $\spadesuit$ 0.9728 $\bullet$ 0.9759 $\bullet$ 0.9770 $\circ$ 0.9746 $\bullet$ 0.9756 $\bullet$ 0.9759 $\bullet$ 0.9770 $\circ$ 0.9766 -0.0004
Yeast_spo5 0.9736 $\bullet$ 0.9711 $\bullet$ 0.9737 $\bullet$ 0.9742 $\spadesuit$ 0.9690 $\bullet$ 0.9698 $\bullet$ 0.9640 $\bullet$ 0.9742 $\spadesuit$ 0.9744 +0.0002
Yeast_spoem 0.9783 $\bullet$ 0.9764 $\bullet$ 0.9789 $\spadesuit$ 0.9790 $\spadesuit$ 0.9750 $\bullet$ 0.9694 $\bullet$ 0.8925 $\bullet$ 0.9790 $\spadesuit$ 0.9792 +0.0001
Table 2: The comparison of label distribution prediction performance measured by Cosine coefficient $\uparrow$.
Evaluation Metrics
Following (Geng, 2016), we adopt six metrics to evaluate label distribution prediction performance: Chebyshev distance (Cheby), Clark distance (Clark), Canberra metric (Canb), Kullback-Leibler divergence (KL), Cosine coefficient (Cos) and Intersection (Inter). By denoting $\bm{\ell}$ and $\hat{\bm{\ell}}$ as ground-truth and predicted label distribution vectors respectively, these metrics are defined as follows:
  • $\mathrm{Cheby}(\bm{\ell},\hat{\bm{\ell}})=\|\bm{\ell}-\hat{\bm{\ell}}\|_{\infty}\downarrow$.
  • $\mathrm{Clark}(\bm{\ell},\hat{\bm{\ell}})=\Big{\|}\frac{\bm{\ell}-\hat{\bm{\ell}}}{\bm{\ell}+\hat{\bm{\ell}}}\Big{\|}_{2}\downarrow$.
  • $\mathrm{Canb}(\bm{\ell},\hat{\bm{\ell}})=\Big{\|}\frac{\bm{\ell}-\hat{\bm{\ell}}}{\bm{\ell}+\hat{\bm{\ell}}}\Big{\|}_{1}\downarrow$.
  • $\mathrm{KL}(\bm{\ell},\hat{\bm{\ell}})=\sum_{y\in\mathcal{Y}}\bm{\ell}^{y}\log\frac{\bm{\ell}^{y}}{\hat{\bm{\ell}}^{y}}\downarrow.$
  • $\mathrm{Cos}(\bm{\ell},\hat{\bm{\ell}})=\frac{\bm{\ell}^{\top}\hat{\bm{\ell}}}{\|\bm{\ell}\|_{2}\|\hat{\bm{\ell}}\|_{2}}\uparrow.$
  • $\mathrm{Inter}(\bm{\ell},\hat{\bm{\ell}})=\sum_{y\in\mathcal{Y}}\min(\bm{\ell}^{y},\hat{\bm{\ell}}^{y})\uparrow.$
Label Distribution Prediction
Table 2 shows the label distribution prediction performance measured by Cosine coefficient, as the averaged scores on ten-fold cross-validation. On each dataset, the best performer and runner-up are respectively highlighted by boldface and underline. We perform paired t tests between the proposed SNEFY-LDL and every baseline method, and use symbols $\bullet$, $\circ$ and $\spadesuit$ to respectively denote that SNEFY-LDL is superior, inferior and equivalent to the baseline at 0.05 significance level. The last column $\Delta$ gives the performance difference between SNEFY-LDL and the best baseline.
From Table 2, we can find that the proposed SNEFY-LDL achieves very competitive label distribution prediction performance, compared with the state-of-the-art baselines. In the majority of cases, the performance of SNEFY-LDL is no worse than the compared baselines. On four datasets, SNEFY-LDL achieves the best Cosine coefficient scores. While SNEFY-LDL does not perform best on the remaining datasets, its performance gap with regard to the best performer is very marginal and even negligible. The performance scores measured by the remaining five metrics are provided in the Appendix, which show the same trend. The competitive label distribution prediction performance implies that the proposed SNEFY-LDL model is able to accurately model the distribution of label distribution vectors, offering an effective tool for label distribution prediction.

### v2

@@ 5 SNEFY-LDL
The proof is provided in the Appendix. With the closed-form kernel function in Eq. (9), we can construct the conditional SNEFY distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ in the form of Eq. (6). This model provides us with the freedom to choose any sufficient statistic $\bm{t}_{2}(\cdot)$ that is used to transform the input sample $\bm{x}$ from the original $d$-dimensional space to the latent $D_{2}$-dimensional space. To capture the non-linearity between input samples and their label distributions, deep neural networks can be leveraged to construct $\bm{t}_{2}(\cdot)$. The input can also be extended beyond the vector-format samples to data with special structures, like images, texts and graphs, where we can respectively leverage Convolutional Neural Networks (CNNs) [Venkatesan and Li, 2017], Transformers [Vaswani et al., 2017] and Graph Neural Networks (GNNs) [Kipf and Welling, 2017] to construct $\bm{t}_{2}(\cdot)$ for end-to-end learning.
The conditional distribution formulation $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ also provides a closed form of mean, variance and co-variance of each label’s composition ratio in describing the conditioning sample $\bm{x}$. About this, we have the following theorem:
@@
Assuming the label distribution vector $\bm{\ell}$ follows the SNEFY conditional distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ in Eq. (6) with the kernel function $\bm{k}_{\sigma,\bm{t}_{1},\bm{t}_{2},\mu_{1}}(\bm{\theta}_{i},\bm{\theta}_{j};\bm{x})$ given in Eq. (9), under the setting that $\bm{t}_{1}(\bm{\ell})=(\log\bm{\ell}^{y_{1}},\log\bm{\ell}^{y_{2}},\cdots,\log\bm{\ell}^{y_{L}})$, $\sigma=\exp$, and $\mu_{1}(d\bm{\ell})=d\bm{\ell}$, as well as the constraint that $\bm{W}_{1}>-1/2$ elementwise, for the $r$th label’s composition ratio, $\bm{\ell}^{y_{r}}$, we have its conditional mean $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ as
@@
The proof is provided in the Appendix. Given the fitted distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$, the mean $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ can be used to predict the unknown label distribution as the expectation over all values in the simplex. We can use the variance $\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]$ to quantify label distribution prediction uncertainties. We can also use $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ with $\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]$ to construct confidence intervals for label distribution predictions by applying Chebyshev’s inequality [Grimmett and Stirzaker, 2020]. The covariance $\mathrm{Cov}[\bm{\ell}^{y_{r}},\bm{\ell}^{y_{s}}|\bm{x}]$ is helpful for us to understand the correlations between two different labels. More importantly, all the statistics are conditioned on the given sample $\bm{x}$, guiding us to make instance-wise decisions.
@@
There are numerous metrics to measure the consistency between two label distributions, like Chebyshev distance, Kullback-Leibler divergence and Cosine coefficient [Geng, 2016]. Instead of optimizing these metrics, the MLE based learning objective in Eq. (16) provides an alternative way to train the LDL model. The fitted distribution $\mathrm{P}(d\bm{\ell}|\bm{x};\bm{V},\bm{\Theta})$ can be applied to various downstream tasks for quantifying the uncertainty of label distribution predictions.
Algorithm Description and Time Complexity. We train the SNEFY-LDL model with stochastic gradient descent. The training procedure is shown in Algorithm 1. The model parameters $\bm{V}$, $\bm{\Theta}$ and $\bm{t}_{2}$ are first initialized by random numbers. We then iteratively select a batch of training samples, calculate the batched likelihoods with Eq. (6), and update parameters $\bm{V}$, $\bm{\Theta}$ and $\bm{t}_{2}$ by descending them along the gradient of batched negative log likelihoods. Taking the epoch number as a constant and assuming the latent layers of $\bm{t}_{2}$ have neurons in the same scale as the neuron number in the last layer $D_{2}$, the time complexity of Algorithm 1 is $\mathcal{O}(N(mn^{2}+Ln^{2}+dD_{2}+D_{2}^{2}))$, which is linear to the number of training samples $N$, making the algorithm able to scale to large datasets. For any sample $\bm{x}$ with an unknown label distribution, its label distribution mean and variance can be computed in time complexity $\mathcal{O}(mn^{2}+Ln^{2}+dD_{2}+D_{2}^{2})$ using the closed-form formulations in Eq. (10) and Eq. (12).
@@
Dataset #Examples #Features #Labels
Movie 7,755 1,869 5
Natural Scene 2,000 294 9
SBU_3DFE 2,500 243 6
SJAFFE 213 243 6
Table 1: Summary of the four benchmark datasets.
@@ 6 Experiments
We conduct extensive experiments on conformal prediction, active learning and ensemble learning to verify SNEFY-LDL’s ability in LDL uncertainty quantification.
Class Id bin size $=2$ bin size $=4$ bin size $=8$
Dirichlet SNEFY-LDL Dirichlet SNEFY-LDL Dirichlet SNEFY-LDL
1 0.8577$\pm$0.0310 0.8747$\pm$0.0243 0.6987$\pm$0.2685 0.8524$\pm$0.0298 0.6476$\pm$0.2505 0.8335$\pm$0.0318
2 0.8512$\pm$0.0333 0.8867$\pm$0.0300 0.4878$\pm$0.3210 0.8723$\pm$0.0361 0.4581$\pm$0.2952 0.8277$\pm$0.0472
3 0.8921$\pm$0.0208 0.8924$\pm$0.0183 0.1644$\pm$0.3396 0.8104$\pm$0.0576 0.1395$\pm$0.2904 0.7537$\pm$0.0788
4 0.8964$\pm$0.0176 0.8804$\pm$0.0256 0.5403$\pm$0.3666 0.8801$\pm$0.0257 0.4897$\pm$0.3246 0.8188$\pm$0.0412
5 0.8492$\pm$0.0300 0.8690$\pm$0.0344 0.7561$\pm$0.2065 0.5276$\pm$0.3452 0.6676$\pm$0.1905 0.4883$\pm$0.3184
6 0.8806$\pm$0.0248 0.8890$\pm$0.0230 0.2005$\pm$0.3537 0.8323$\pm$0.1615 0.1817$\pm$0.3193 0.8149$\pm$0.1582
7 0.8447$\pm$0.0335 0.8662$\pm$0.0303 0.5131$\pm$0.3417 0.8499$\pm$0.0340 0.4827$\pm$0.3210 0.7697$\pm$0.0754
8 0.8427$\pm$0.0329 0.8897$\pm$0.0218 0.5107$\pm$0.3399 0.8893$\pm$0.0219 0.4840$\pm$0.3175 0.8753$\pm$0.0245
9 0.8107$\pm$0.0548 0.7681$\pm$0.0472 0.1443$\pm$0.2992 0.4513$\pm$0.2982 0.1242$\pm$0.2609 0.4096$\pm$0.2741
Table 2: The conformal prediction performance measured by the FSC metric on the Natural_Scene dataset.
Class Id bin size $=2$ bin size $=4$ bin size $=8$
Dirichlet SNEFY-LDL Dirichlet SNEFY-LDL Dirichlet SNEFY-LDL
1 0.8913$\pm$0.0205 0.8948$\pm$0.0175 0.8407$\pm$0.0511 0.8836$\pm$0.0189 0.5391$\pm$0.3022 0.8277$\pm$0.0675
2 0.8816$\pm$0.0248 0.8658$\pm$0.0297 0.8703$\pm$0.0266 0.8472$\pm$0.0364 0.8475$\pm$0.0326 0.8210$\pm$0.0435
3 0.8737$\pm$0.0278 0.8963$\pm$0.0175 0.8552$\pm$0.0292 0.8785$\pm$0.0258 0.8290$\pm$0.0349 0.8505$\pm$0.0297
4 0.8848$\pm$0.0238 0.8879$\pm$0.0222 0.8274$\pm$0.0367 0.8725$\pm$0.0270 0.7193$\pm$0.1287 0.7261$\pm$0.1500
5 0.8937$\pm$0.0177 0.8955$\pm$0.0184 0.8150$\pm$0.0462 0.8381$\pm$0.0403 0.0793$\pm$0.2388 0.4199$\pm$0.3376
6 0.8911$\pm$0.0191 0.8920$\pm$0.0205 0.8625$\pm$0.0366 0.8334$\pm$0.0453 0.6025$\pm$0.2969 0.5048$\pm$0.3136
Table 3: The conformal prediction performance measured by the FSC metric on the SBU_3DFE dataset.
@@
Benchmark Datasets. We use four datasets [Geng, 2016] to benchmark our experiments, including the Movie dataset containing label distributions on five movie rating scales, the Natural Scene dataset with label distributions constructed by inconsistent multi-label ranking on natural scene images, the facial expression datasets SBU_3DFE and SJAFFE with label distributions on six emotions. The statistics of the four benchmark datasets are summarized in Table 1.
@@
Implementation Details. When implementing SNEFY-LDL, $n$ and $m$ are respectively set to 64 and 32, $D_{2}$ is set as equal to $n$ and a one-layer neural network with ReLU activation is used to construct $\bm{t}_{2}$. The model is trained for 100 epochs with batch size 64 for conformal prediction and batch size 16 for active learning and ensemble learning. Weight clipping [Arjovsky et al., 2017] is used to control $\bm{W}_{1}>-1/2$ after each parameter update.
6.1 Conformal Prediction
As shown in Theorem 3, with the trained SNEFY-LDL model, given a new sample $\bm{x}$, we can get the closed-form conditional mean for the $r$th label’s composition ratio as $\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]$ in Eq. (10) and the closed-form conditional variance $\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]$ in Eq. (12). By applying the Chebyshev’s inequality [Grimmett and Stirzaker, 2020], we can construct a confidence interval for the $r$th label’s composition ratio in describing $\bm{x}$, $\bm{\ell}^{y_{r}}_{\bm{x}}$, which can be formally stated as
$\mathrm{P}(|\bm{\ell}^{y_{r}}_{\bm{x}}-\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]|\leq k\sqrt{\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]})\geq 1-\frac{1}{k^{2}},$ (17)
i.e., the confidence interval of $\bm{\ell}^{y_{r}}_{\bm{x}}$ at the $1-1/k^{2}$ level is
$\left[\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]-k\sqrt{\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]},\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]+k\sqrt{\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]}\right],$
which can be further calibrated as a conformal prediction task for the one-dimensional uncertainty estimate $\bm{\ell}^{y_{r}}_{\bm{x}}$. According to Angelopoulos and Bates [2021], on a calibration set with $N_{\text{cal}}$ samples, we can define a calibration score function as
$s(\bm{x},\bm{\ell}^{y_{r}}_{\bm{x}})=\frac{\lvert\bm{\ell}^{y_{r}}_{\bm{x}}-\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]\rvert}{k\sqrt{\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]}},$ (18)
where $\bm{\ell}^{y_{r}}_{\bm{x}}$ is the ground-truth value of the $r$th label’s composition ratio in describing the calibration sample $\bm{x}$. By computing the calibration scores of all calibration samples, we can get the $\lceil(1-1/k^{2})(N_{\text{cal}}+1)\rceil/N_{\text{cal}}$ quantile of the calibration scores as $\hat{q}_{y_{r}}$. For a new sample $\bm{x}$, the calibrated confidence interval for $\bm{\ell}^{y_{r}}_{\bm{x}}$ at the $1-1/k^{2}$ level is
$\mathcal{C}(\bm{x},\bm{\ell}^{y_{r}})=\left[\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]-k\cdot\hat{q}_{y_{r}}\sqrt{\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]},\mathrm{E}[\bm{\ell}^{y_{r}}|\bm{x}]+k\cdot\hat{q}_{y_{r}}\sqrt{\mathrm{Var}[\bm{\ell}^{y_{r}}|\bm{x}]}\right].$ (19)